# Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits

### v2

@@ Appendix E Missing Proof from Appendix D
$\leq\frac{1}{2S}(\exp(2\delta)-1)^{2}\leq\frac{36\delta^{2}}{2S}=\frac{18}{n},$ (E.19)
where the last inequality follows from $\exp(x)-1\leq 3x$ for $x\in[0,0.5]$ and $\delta=\sqrt{S/n}\leq 0.25$ by assumption. Therefore, we substitute E.18 into Assouad’s Lemma (Lemma F.3) to obtain
LHS of E.15 $\geq\frac{1}{S}\cdot S\cdot\frac{\eta S}{\alpha n}\cdot\frac{1}{4}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big{)}$
$=0.25\cdot\frac{\eta S}{\alpha n}\cdot\exp\Big{(}-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big{)}\geq\frac{\eta S}{\alpha n}\cdot\frac{1}{3}\cdot\exp(-18)\gtrsim\frac{\eta S}{\alpha n},$ (E.20)
where the $1/S$ comes from the denominator of E.16 and the second inequality follows from E.19. ∎
Appendix F Technical Results and Standard Results
Lemma F.1 (Zhao et al. 2024, Lemma D.4).
Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 4.1. Suppose ${\bar{g}}$ is the result of MLE estimation of Algorithm 3, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
@@ Appendix F Technical Results and Standard Results
Lemmas F.2 and F.3 are two standard reductions (Le Cam, 1973; Yu, 1997; Polyanskiy and Wu, 2025). See, e.g., Chen et al. (2024, Section 3) for a general proof.
Lemma F.2 (Fano’s inequality).
Fix any :-$\mathcal{R}\coloneq\{r_{1},\cdots,r_{S}\}$ and policy class $\Pi$, let $L:\Pi\times\mathcal{R}\to\mathbb{R}_{+}$ be some loss function. Suppose there exist some constant $c>0$ such that the following condition holds:
@@
$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq\frac{c}{2}\bigg{(}1-\frac{\max_{i\neq j}\mathsf{KL}(P_{r_{i}}\|P_{r_{j}})+\log 2}{\log S}\bigg{)},$
where $P_{r}$ is the distribution of dataset given model $r\in\mathcal{R}$.
Lemma F.3 (Assouad’s Lemma).
@@
$\theta\sim_{j}\theta^{\prime}\Rightarrow\inf_{\pi\in\Pi}L_{j}(\pi,r_{\theta})+L_{j}(\pi,r_{\theta^{\prime}})\geq c$ (F.1)
@@
$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq S\cdot\frac{c}{4}\min_{\exists j:\theta\sim_{j}\theta^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{r_{\theta}}\|P_{r_{\theta^{\prime}}}\right)\Big{)},$
where $P_{r}$ denotes the distribution of the dataset given $r\in\mathcal{R}$.
The following Lemma F.4 is due to Gilbert (1952); Varshamov (1957), which is a classical result in coding theory.
Lemma F.4.

### v3

@@ Appendix G Missing proof from Appendix F
$\leq\frac{1}{2S}(\exp(2\delta)-1)^{2}\leq\frac{36\delta^{2}}{2S}=\frac{18}{n},$ (G.20)
where the last inequality follows from $\exp(x)-1\leq 3x$ for $x\in[0,0.5]$ and $\delta=\sqrt{S/n}\leq 0.25$ by assumption. Therefore, we substitute G.19 into Assouad’s Lemma (Lemma H.3) to obtain
LHS of G.16 $\geq\frac{1}{S}\cdot S\cdot\frac{\eta S}{\alpha n}\cdot\frac{1}{4}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big(-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big)$
$=0.25\cdot\frac{\eta S}{\alpha n}\cdot\exp\Big(-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big)\geq\frac{\eta S}{\alpha n}\cdot\frac{1}{3}\cdot\exp(-18)\gtrsim\frac{\eta S}{\alpha n},$ (G.21)
where the $1/S$ comes from the denominator of G.17 and the second inequality follows from G.20. ∎
Appendix H Auxiliary Lemmas
Lemma H.1 (Zhao et al. 2024, Lemma D.4).
Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section F.1. Suppose $\bar{g}$} is the result of MLE estimation of Algorithm 3, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big[{\bar{g}}(s,a)-{g^{*}}(s,a)\big]$, then with probability at least $1-2\delta$, we have
@@ Appendix H Auxiliary Lemmas
Lemmas H.2 and H.3 are two standard reductions (Le Cam, 1973; Yu, 1997; Polyanskiy and Wu, 2025). See, e.g., Chen et al. (2024, Section 3) for a general proof.
Lemma H.2 (Fano’s inequality).
Fix any $\mathcal{R}\coloneq\{r_{1},\cdots,r_{S}\}$ and policy class $\Pi$, let $L:\Pi\times\mathcal{R}\to\mathbb{R}_{+}$ be some loss function. Suppose there exist some constant $c>0$ such that the following condition holds:
@@
$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}\mathbb{E}_{\mathcal{D}\sim P_{r}}L\big(\pi(\mathcal{D}),r\big)\geq\frac{c}{2}\bigg(1-\frac{\max_{i\neq j}\mathsf{KL}(P_{r_{i}}\|P_{r_{j}})+\log 2}{\log S}\bigg),$
where the trajectory distribution of $\pi$ interacting with instance $r\in\mathcal{R}$ is denoted by $P_{r}$.
Lemma H.3 (Assouad’s Lemma).
@@
$\theta\sim_{j}\theta^{\prime}\Rightarrow\inf_{\pi\in\Pi}L_{j}(\pi,r_{\theta})+L_{j}(\pi,r_{\theta^{\prime}})\geq c$ (H.1)
@@
$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}\mathbb{E}_{\mathcal{D}\sim P_{r}}L\big(\pi(\mathcal{D}),r\big)\geq S\cdot\frac{c}{4}\min_{\exists j:\theta\sim_{j}\theta^{\prime}}\exp\Big(-{\mathsf{KL}}\left(P_{r_{\theta}}\|P_{r_{\theta^{\prime}}}\right)\Big),$
where the trajectory distribution of $\pi$ interacting with instance $r\in\mathcal{R}$ is denoted by $P_{r}$.
The following Lemma H.4 is due to Gilbert (1952); Varshamov (1957), which is a classical result in coding theory.
Lemma H.4.